OpenStudy (anonymous):

Do i use the foil solution for? g^2 -9

OpenStudy (anonymous):

You can also try this formula: \[\Large a^{2}-b^{2}=(a-b)(a+b)\]

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

g^2 -9^2 =( g -9) (g+9)

OpenStudy (anonymous):

Well 9 is really (3)^2

OpenStudy (anonymous):

Okay brb give me a few minutes getting ready for bed

OpenStudy (anonymous):

Back, thanks for waiting.

OpenStudy (anonymous):

So g^2 -3^2 = (g - 3)(g +3) Is this right?

OpenStudy (anonymous):

@doulikepiecauseidont Are you still there?

jimthompson5910 (jim_thompson5910):

(g - 3)(g +3) is correct

OpenStudy (anonymous):

Thanks Though i wasn't going to get that one finished today

OpenStudy (anonymous):

Yes it's right